OpenStudy (anonymous):

What is the simplest form of the expression? sqrt 20 +sqrt 45 -sqrt5 4sqrt5 6sqrt5 13sqrt5 5sqrt5

OpenStudy (binary3i):

\[\sqrt{20} + \sqrt{45} -\sqrt{5}=\sqrt{5*4} +\sqrt{5*9} -\sqrt{5}\] \[\sqrt{5} (\sqrt{4} +\sqrt{9} -1)= \sqrt{5}(2+3-1)=4\sqrt{5}\]